Could someone help please!
1.) If x^2-16=3 then x must approximately be ? (This is all under the radial sign)
2.) If (5x-2)^2=6 then the solution set for x must be ?
3.) What expression is equivalent to 27^-3/2
4.) What is the domain of the function f(x)=1/(x-1)^2+3

Answer by stanbon(75887)

1.) If x^2-16=3 then x must approximately be ? (This is all under the radial sign)
What is under the radical?
---------------------
2.) If (5x-2)^2=6 then the solution set for x must be ?
Take the square root of both sides to get:
5x-2=sqrt6 or 5x-2=-sqrt6
x=(2+sqrt6)/5 or x=(2-sqrt6)/5
--------------
3.) What expression is equivalent to 27^-3/2
Comment: -3/2 is the exponent.
the - means "invert"
the 3 means "cube"
the 2 means "sqrt root"
You may do these in any order.
-1/sqrt(27^3)
4.) What is the domain of the function f(x)=1/(x-1)^2+3
Since (x-1) is in the denominator, x cannot be 1.
Domain: All real numbers except x=1.
Cheers,
Stan H.